### v1

@@ Abstract
Federated learning is a machine learning method that supports training models on decentralized devices or servers, where each holds its local data, removing the need for data exchange. This approach is especially useful in healthcare, as it enables training on sensitive data without needing to share them. The nature of federated learning necessitates robust security precautions due to data leakage concerns during communication. To address this issue, we propose a new approach that employs selective encryption, homomorphic encryption, differential privacy, and bit-wise scrambling to minimize data leakage while achieving good execution performance. Our technique , FAS (fast and secure federated learning) is used to train deep learning models on medical imaging data. We implemented our technique using the Flower framework and compared with a state-of-the-art federated learning approach that also uses selective homomorphic encryption. Our experiments were run in a cluster of eleven physical machines to create a real-world federated learning scenario on different datasets. We observed that our approach is up to 90% faster than applying fully homomorphic encryption on the model weights. In addition, we can avoid the pretraining step that is required by our competitor and can save up to 20% in terms of total execution time. While our approach was faster, it obtained similar security results as the competitor.
Keywords:
@@
1 Introduction
Federated learning, a technique introduced by Google in 2016, offers a way to train machine learning models across decentralized data from mobile devices (federated, ). This approach has since gained widespread interest in both academic research and industry applications. By allowing model training on data that remains locally stored, federated learning helps protect privacy, making it an attractive solution for sensitive data environments. For example, in healthcare, it enables institutions to build shared machine learning models without the need to transfer patient data, thereby maintaining compliance with data protection laws. Overall, federated learning facilitates secure, privacy-conscious model training across decentralized networks.
@@ 1 Introduction
The rest of this paper is organized as follows: The Datasets section describes the different image datasets used in our study. In Background on Deep Learning Models and Security Techniques, we discuss the deep learning models evaluated in our experiments, including ResNet-50 (resnet, ), DenseNet121 (densenet, ), EfficientNetB0 (efficientnet, ), MobileNet V2 (mobilenet, ), and elaborate on the security techniques applied.
The Experimental Setup for Federated Learning details our use of CloudLab (cloudlab, ) as the testbed and the Flower framework (flower, ) for federated learning experiments. The Evaluation section presents encryption overhead and time costs across datasets and models, with a detailed time comparison of security techniques.
In Comparison of Security Techniques, we evaluate differential privacy (differential, ), homomorphic encryption (gentry2009, ), and FAS, and compare these with similar works. The Practical Implications section discusses efficiency considerations and the balance between security and performance, followed by a Visual Comparison of FAS Approaches. Finally, we conclude with a summary of our findings and future research directions.
2 Background and Related work
Privacy-preserving techniques in federated learning have garnered significant attention due to the increasing need to protect sensitive data during model training across distributed devices. Federated learning, introduced by McMahan et al. (mcmahan2017communication, ), enables the training of models on decentralized data sources, but it also introduces vulnerabilities that could potentially leak private information through model updates.
Homomorphic encryption (HE) has emerged as a robust method for maintaining data confidentiality in federated learning. It allows computations to be performed directly on encrypted data, ensuring that sensitive information remains protected throughout the training process. Gentry’s foundational work on fully homomorphic encryption (FHE) (gentry2009, ) made arbitrary computations on encrypted data feasible. However, the high computational and storage overhead of FHE has limited its practical application, particularly in resource-constrained environments (boneh2006, ; brakerski2014efficient, ).
2.1 Efficiency Enhancements in Homomorphic Encryption
@@
Fan and Vercauteren’s BGV Scheme (fan2012somewhat, ): This scheme optimizes the bootstrapping process, making HE more practical by efficiently handling both addition and multiplication on ciphertexts. This approach has been critical in making HE viable for more complex computations in federated learning environments (smart2014fully, ). Xu et al. xu2019hybridalpha proposed HybridAlpha, an efficient privacy-preserving federated learning system that optimizes both privacy and performance. In comparison to our FAS approach, HybridAlpha leverages different privacy mechanisms such as differential privacy and secure aggregation to ensure efficient federated learning while maintaining data confidentiality.
@@
Chillotti et al.’s TFHE Scheme (chillotti2016faster, ): TFHE introduced a faster bootstrapping process, making it highly efficient for secure Boolean computations, thus expanding the potential use of FHE in time-sensitive environments (chillotti2018faster, ).
@@
Ducas and Micciancio’s FHEW Scheme (ducas2015faster, ): FHEW further improved the speed of homomorphic operations by optimizing the bootstrapping process, making FHE more feasible for larger-scale applications (ducas2015faster, ).
@@
Cheon et al.’s CKKS Scheme (cheon2017homomorphic, ): This scheme supports approximate arithmetic operations on encrypted data, making it particularly efficient for machine learning and statistical analysis tasks where exact precision is not always required (cheon2018bootstrapping, ).
@@
Brakerski, Gentry, and Vaikuntanathan’s BGV+ Scheme (brakerski2014efficient, ): This work built on the BGV scheme, introducing further optimizations that reduce the computational overhead of FHE, particularly in noise management, which is crucial for maintaining the accuracy of federated learning models under encryption (brakerski2014efficient, ).
2.2 Differential Privacy (DP) in Federated Learning
Differential privacy (DP) introduces controlled noise into the data or model updates, protecting individual data points while allowing accurate aggregate analysis. The foundational work by Dwork et al. (dwork2006calibrating, ) laid the groundwork for DP, and its application in deep learning was further developed by Abadi et al. (abadi2016deep, ), who integrated DP into the stochastic gradient descent (SGD) process. Subsequent research has focused on balancing the trade-off between privacy guarantees and model accuracy, with methods such as the moment accountant providing more precise privacy loss tracking (mironov2017renyi, ).
2.3 Selective Encryption in Federated Learning
Selective encryption provides a middle ground by encrypting only a subset of the model parameters, thereby reducing the computational burden while maintaining a reasonable level of security. Wu et al. (wu2019towards, ) demonstrated that selective encryption could significantly reduce overhead without compromising overall security by carefully choosing which parameters to encrypt. Other studies have explored similar strategies, focusing on optimizing the trade-off between encryption and efficiency in federated learning systems (li2019privacy, ; song2020privacy, ).
In this paper, we extend these concepts by introducing a novel selective parameter encryption scheme that incorporates differential noise and bitwise scrambling. This approach not only reduces the overhead associated with homomorphic encryption but also enhances privacy protection for the unencrypted parameters. By leveraging the efficiency of selective encryption and the principles of differential privacy, our method provides a practical solution for secure and efficient federated learning in sensitive domains.
2.4 Bitwise Scrambling in Federated Learning: Enhancing Security with Low Overhead
Bitwise scrambling is an obfuscation technique that further protects partly encrypted data by rearranging bits according to a cryptographic key, making it difficult for adversaries to extract meaningful information without access to the key. While scrambling techniques have long been used in secure communication protocols, recent research has adapted bitwise scrambling specifically for federated learning environments to enhance privacy with minimal computational overhead.
Yang et al. (yang2017bitwise, ) explored the use of bitwise scrambling as an additional layer of security in data transmission. By applying scrambling to unencrypted or selectively encrypted data, this technique provides a cost-effective way to obfuscate sensitive information during federated model training and aggregation. In federated learning, bitwise scrambling effectively prevents adversaries from performing direct analysis or reconstructing original data from partly encrypted models, even if the differential privacy noise is circumvented.
Other studies have further demonstrated the value of combining bitwise scrambling with homomorphic encryption in federated learning frameworks. For example, Halevi et al. (halevi2019faster, ) integrated scrambling with HE to enable secure, efficient transmission in distributed computing, showing that scrambling can effectively prevent chosen-ciphertext attacks by masking encrypted data patterns. This approach not only reduces the attack surface but also allows for faster processing times, as only a portion of the data requires full encryption while scrambled data remains obfuscated.
Bitwise scrambling, when combined with selective encryption and differential privacy as presented in this paper, enables a highly efficient privacy-preserving solution for federated learning. It obfuscates sensitive information without introducing significant additional computation, making it suitable for resource-constrained environments that still require robust security protections.
3 Motivation and Threat Model
3.1 Motivation
@@
The growing adoption of federated learning (FL) in sensitive domains, such as healthcare and finance, has heightened the need for robust privacy-preserving mechanisms. While federated learning inherently minimizes the risk of data exposure by keeping raw data localized, it is still susceptible to a variety of privacy threats due to the exchange of model updates during training. Attackers can exploit these updates to infer sensitive information, such as patient health records or financial transactions, through techniques like model inversion or gradient leakage. These vulnerabilities necessitate enhanced security measures that preserve privacy without compromising performance.
@@
Existing privacy-preserving techniques, such as full homomorphic encryption and differential privacy, have demonstrated effectiveness but face practical challenges. Fully homomorphic encryption ensures the highest level of data confidentiality but introduces significant computational and communication overhead, making it unsuitable for resource-constrained or time-sensitive environments. Differential privacy, while lightweight, can lead to degradation in model accuracy due to the introduction of noise, particularly in scenarios requiring high precision.
@@
To address these challenges, this work introduces a novel FAS approach that balances security and efficiency. By selectively encrypting critical portions of the model updates and applying differential noise and bitwise scrambling to the remaining data, FAS reduces computational overhead while maintaining strong privacy guarantees. This method is especially valuable in federated learning scenarios where computational resources and latency are critical, such as medical imaging and real-time analytics.
The proposed approach provides a scalable, efficient, and robust solution to protect sensitive data in FL, bridging the gap between the security of full homomorphic encryption and the efficiency of lightweight techniques like differential privacy.
3.2 Threat Model
@@
Federated learning (FL) systems are vulnerable to privacy threats due to the decentralized nature of training and the exchange of model updates. Adversaries can exploit these updates to reconstruct sensitive training data, such as medical images, or infer private attributes through techniques like gradient inversion and statistical analysis. This threat model assumes the presence of a semi-honest adversary who follows the protocol but attempts to infer sensitive information from intercepted or compromised data.
@@
The adversary is capable of accessing model updates exchanged between clients and the central server and may compromise either the server or individual clients to gain insights. Potential attack vectors include gradient inversion attacks to reconstruct training data, membership inference to determine if specific data points were included in the training, and statistical analysis to extract private attributes. While all communications are assumed to be encrypted (e.g., using SSL), this does not prevent adversaries from analyzing the transmitted updates.
@@
To mitigate these threats, our approach combines selective encryption, differential noise addition, and bitwise scrambling. Critical model parameters are encrypted, while unencrypted data is obfuscated with noise and scrambling to prevent reconstruction. Additionally, secure aggregation ensures that data remains protected throughout the learning process, providing robust defenses against privacy attacks without incurring excessive computational overhead. This layered strategy strikes a balance between security and efficiency, making it well-suited for sensitive applications like healthcare.
4 Proposed Model
@@ 4.1 Selective Encryption
4.2 Differential Noise Addition
@@ 4.3 Bitwise Scrambling
4.4 Federated Learning Integration
@@
4.5 Security and Efficiency Trade-Off
@@
4.6 Privacy-Preserving Techniques
4.6.1 Homomorphic Encryption
@@
where Enc and Dec denote encryption and decryption, respectively, and Eval represents a computation on encrypted data. This property is critical for privacy-preserving applications, as it prevents adversaries from accessing plaintext data (gentry2009, ).
4.6.2 Differential Privacy
@@
where $\epsilon$ is the privacy budget, controlling the trade-off between privacy and utility (dwork2006calibrating, ).
@@
4.7 Process Flow of the Proposed Model
@@
4.8 Federated Learning System Process Flow
@@
4.8.1 Client-Side Processing
@@
4.8.2 Server-Side Processing
@@
4.8.3 Client Post-Processing
@@
4.9 Advantages of the Proposed Model
@@ 4.10.1 Setup and Definitions
Parameters in $S$ are encrypted using a semantically secure HE scheme. Semantic security ensures no polynomial-time adversary can distinguish ciphertexts of different messages, implying no additional privacy cost in a DP sense (0-DP).
@@
For each $i\in[N]\setminus S$, we add noise calibrated to $\epsilon_{i}$-DP. By the composition property of DP, releasing all these parameters together is $\left(\sum_{i\in[N]\setminus S}\epsilon_{i}\right)$-DP.
@@
We define a scrambling function $\mathcal{T}_{k}:\mathcal{X}\to\mathcal{X}$, where $\mathcal{X}$ is the space of possible model outputs. $\mathcal{T}_{k}$ is deterministic, keyed by $k$, and independent of $D$ except through $M(D)$. This scrambling is a form of post-processing.
4.10.2 Privacy Guarantee
Theorem 4.1
@@ 4.10.2 Privacy Guarantee
    Encrypts $W_{i}$ for $i\in S$ using a semantically secure HE scheme (gentry2009, ; Fontaine2007, ).
@@
    Adds noise to each $W_{j}$ for $j\in[N]\setminus S$ to achieve $\epsilon_{j}$-DP. Releasing all noisy parameters together is $\left(\sum_{j\in[N]\setminus S}\epsilon_{j}\right)$-DP (Dwork2014, ).
@@
    Applies a scrambling function $\mathcal{T}_{k}$ to the entire output, resulting in $\mathcal{M}^{\prime}(D)=\mathcal{T}_{k}(\mathcal{M}(D))$ (Dwork2006, ).
@@
Proof
Step 1 (DP on $[N]\setminus S$): For $[N]\setminus S$, each parameter $W_{j}+\text{noise}_{j}$ is $\epsilon_{j}$-DP. By composition, releasing all these parameters is $\sum_{j\in[N]\setminus S}\epsilon_{j}$-DP (Dwork2014, ). Formally, for any adjacent $D,D^{\prime}$ and event $O$:
@@
Step 2 (Selective Encryption on $S$): Parameters in $S$ are encrypted with a semantically secure HE scheme. Without the secret key, no information about the plaintext is revealed, contributing 0-DP cost (gentry2009, ; Fontaine2007, ).
Step 3 (Scrambling as Post-Processing): The scrambling $\mathcal{T}_{k}$ is a deterministic, data-independent post-processing step. A fundamental property of DP is closure under post-processing: if $\mathcal{M}$ is $\epsilon$-DP, then $f(\mathcal{M})$ is also $\epsilon$-DP for any deterministic $f$ (Dwork2006, ; Dwork2014, ). Since $\mathcal{T}_{k}$ does not depend on $D$ beyond $\mathcal{M}(D)$ itself, applying $\mathcal{T}_{k}$ cannot increase or decrease the DP budget.
@@
Thus, $\mathcal{M}^{\prime}(D)$ maintains the same DP guarantee as $\mathcal{M}(D)$.
5 Evaluation
In this section, we evaluate our federated learning setup using multiple security techniques applied to different image datasets. These techniques include homomorphic encryption, differential privacy, and our custom FAS. We assessed their effectiveness based on several metrics, such as test accuracy, communication cost, computational overhead, and the overall ability to maintain data privacy and security in a federated learning context.
The datasets were split into 10 equal-sized partitions and distributed among machines running Flower clients. Each machine trained its local model on its respective partition, with training occurring over 10 rounds of 20 epochs each. We used 10% of the training data to construct a validation set, ensuring consistent evaluation throughout the process. After the rounds were completed, the final model accuracy was measured on a separate test set for each dataset. Notably, all deep learning models were trained using only CPU resources to simulate realistic resource constraints.
The following tables compare the training times (in minutes) for different models across various datasets: CIFAR-10, CT Kidney, Lung, and diabetic retinophaty. Each table presents three encryption configurations:
5.1 Datasets used in experiments
To ensure a comprehensive evaluation of our model across various imaging contexts, we selected publicly available datasets with high usability ratings. Our criteria for dataset selection included broad coverage of diseases, utilization of diverse imaging techniques, and a substantial quantity of images for both training and validation purposes. With these requirements in mind, we chose the CIFAR-10 dataset for comparison, along with a medical dataset from Kaggle focusing on Lung Cancer Histopathological images. Below, we provide an overview of each dataset.
5.1.1 CIFAR-10 Dataset
The CIFAR-10 dataset is a widely recognized benchmark in image classification research. It contains 10 classes, each with 6,000 images of 32×32 pixels. The dataset includes 50,000 images for training and 10,000 images for testing, providing a diverse set of images for evaluating model performance. CIFAR-10 is frequently used for benchmarking the effectiveness of various models due to its relatively moderate complexity and diverse set of images (CIFAR10, ).
5.1.2 Chest X-ray (COVID-19, Pneumonia) Dataset
The Chest X-ray dataset includes 6,339 images classified into three categories: benign, COVID-19, and pneumonia, with 2,313 images per category. The grayscale X-ray images provide detailed views of internal body structures, assisting in the assessment of lung conditions. Representative images from each category are shown in Figure 4(ChestXray, )

### v2

@@ Abstract
Federated learning is a machine learning method that supports training models on decentralized devices or servers, where each holds its local data, removing the need for data exchange. This approach is especially useful in healthcare, as it enables training on sensitive data without needing to share them. The nature of federated learning necessitates robust security precautions due to data leakage concerns during communication. To address this issue, we propose a new approach that employs selective encryption, homomorphic encryption, differential privacy, and bit-wise scrambling to minimize data leakage while achieving good execution performance. Our technique FAS (fast and secure federated learning) is used to train deep learning models on medical imaging data. We implemented our technique using the Flower framework and compared with a state-of-the-art federated learning approach that also uses selective homomorphic encryption. Our experiments were run in a cluster of eleven physical machines to create a real-world federated learning scenario on different datasets. We observed that our approach is up to 90% faster than applying fully homomorphic encryption on the model weights. In addition, we can avoid the pretraining step that is required by our competitor and can save up to 46% in terms of total execution time. While our approach was faster, it obtained similar security results as the competitor.
Index Terms:
@@
I Introduction
Federated learning, a technique introduced by Google in 2016, offers a way to train machine learning models across decentralized data from mobile devices [1]. This approach has since gained widespread interest in both academic research and industry applications. By allowing model training on data that remains locally stored, federated learning helps protect privacy, making it an attractive solution for sensitive data environments. For example, in healthcare, it enables institutions to build shared machine learning models without the need to transfer patient data, thereby maintaining compliance with data protection laws. Overall, federated learning facilitates secure, privacy-conscious model training across decentralized networks.
@@ I Introduction
The rest of this paper is organized as follows: The Datasets section describes the different image datasets used in our study. In Background on Deep Learning Models and Security Techniques, we discuss the deep learning models evaluated in our experiments, including ResNet-50 [2], DenseNet121 [3], EfficientNetB0 [4], MobileNet V2 [5], and elaborate on the security techniques applied.
The Experimental Setup for Federated Learning details our use of CloudLab [10] as the testbed and the Flower framework [11] for federated learning experiments. The Evaluation section presents encryption overhead and time costs across datasets and models, with a detailed time comparison of security techniques.
In Comparison of Security Techniques, we evaluate differential privacy [15], homomorphic encryption [16], and FAS, and compare these with similar works. The Practical Implications section discusses efficiency considerations and the balance between security and performance, followed by a Visual Comparison of FAS Approaches. Finally, we conclude with a summary of our findings and future research directions.
II Background and Related Work
Privacy-preserving techniques in federated learning have gained attention due to the need to protect sensitive data across distributed devices. Federated learning, introduced by McMahan et al. [23], enables decentralized model training but poses risks of data leakage through model updates.
Homomorphic encryption (HE) offers a robust solution, allowing computations on encrypted data while maintaining confidentiality. Gentry’s fully homomorphic encryption (FHE) [16] made arbitrary encrypted computations feasible, though its high computational overhead limits practical use [17, 28].
II-A Efficiency Enhancements in Homomorphic Encryption
@@
Fan and Vercauteren’s BGV Scheme [25]: This scheme optimizes the bootstrapping process, making HE more practical by efficiently handling both addition and multiplication on ciphertexts. This approach has been critical in making HE viable for more complex computations in federated learning environments [33]. Xu et al. [41] proposed HybridAlpha, an efficient privacy-preserving federated learning system that optimizes both privacy and performance. In comparison to our FAS approach, HybridAlpha leverages different privacy mechanisms such as differential privacy and secure aggregation to ensure efficient federated learning while maintaining data confidentiality.
@@
Chillotti et al.’s TFHE Scheme [26]: TFHE introduced a faster bootstrapping process, making it highly efficient for secure Boolean computations, thus expanding the potential use of FHE in time-sensitive environments [34].
@@
Ducas and Micciancio’s FHEW Scheme [27]: FHEW further improved the speed of homomorphic operations by optimizing the bootstrapping process, making FHE more feasible for larger-scale applications [27].
@@
Cheon et al.’s CKKS Scheme [29]: This scheme supports approximate arithmetic operations on encrypted data, making it particularly efficient for machine learning and statistical analysis tasks where exact precision is not always required [36].
@@
Brakerski, Gentry, and Vaikuntanathan’s BGV+ Scheme [28]: This work built on the BGV scheme, introducing further optimizations that reduce the computational overhead of FHE, particularly in noise management, which is crucial for maintaining the accuracy of federated learning models under encryption [28].
II-B Differential Privacy (DP) in Federated Learning
Differential privacy (DP) mitigates privacy risks by injecting controlled noise into model updates. Initially proposed by Dwork et al. [30], its integration into deep learning was further developed by Abadi et al. [31]. Subsequent work has refined DP methods to balance privacy and model accuracy [35].
II-C Selective Encryption in Federated Learning
Selective encryption reduces computational burden by encrypting only a portion of model parameters while maintaining security. Wu et al. [32] demonstrated its effectiveness in reducing overhead without compromising privacy. Similar approaches focus on optimizing encryption efficiency in federated learning [37, 38]. Zhang et al. [45] proposed BatchCrypt, an efficient homomorphic encryption framework that batches model updates to reduce the computational and communication overhead in federated learning. While BatchCrypt improves efficiency compared to traditional HE-based methods, it still incurs non-trivial overhead. Our approach further optimizes encryption by selectively encrypting a fraction of updates, achieving similar security guarantees with significantly lower computational costs. This paper extends these concepts by integrating selective encryption with differential noise and bitwise scrambling, reducing homomorphic encryption overhead while enhancing privacy protection.
II-D Bitwise Scrambling for Federated Learning Security
Bitwise scrambling further obfuscates partially encrypted data by rearranging bits using a cryptographic key. This technique complicates adversarial analysis and enhances security without significant computational cost.
Yang et al. [24] applied bitwise scrambling to secure data transmission, effectively preventing direct reconstruction of encrypted models. Halevi et al. [39] demonstrated its use with HE to resist chosen-ciphertext attacks while maintaining efficiency.
Combining selective encryption, differential privacy, and bitwise scrambling offers an effective privacy-preserving solution for federated learning, balancing security and computational efficiency for resource-constrained applications.
III Motivation and Threat Model
III-A Motivation
@@
The adoption of federated learning (FL) in sensitive domains such as healthcare and finance necessitates robust privacy-preserving mechanisms. While FL mitigates data exposure by keeping raw data localized, model updates remain vulnerable to attacks like model inversion and gradient leakage, which can reveal sensitive information such as patient records or financial transactions.
@@
Existing privacy-preserving techniques, such as fully homomorphic encryption (FHE) and differential privacy (DP), have limitations. FHE ensures confidentiality but incurs significant computational and communication overhead, making it impractical for resource-constrained settings. DP, though computationally efficient, can degrade model accuracy by introducing noise, which is problematic in high-precision applications.
@@
To address these challenges, this work proposes the FAS approach, which balances security and efficiency by selectively encrypting critical model updates while applying differential noise and bitwise scrambling to the remaining data. This reduces computational overhead while maintaining strong privacy guarantees, making it particularly suitable for federated learning in time-sensitive applications like medical imaging and real-time analytics.
III-B Threat Model
@@
Federated learning systems are susceptible to privacy threats due to decentralized training and the exchange of model updates. A semi-honest adversary may follow the protocol but attempt to infer sensitive information from intercepted or compromised data.
@@
Adversaries can access model updates between clients and the central server, potentially compromising either to extract insights. Common attack vectors include gradient inversion to reconstruct training data, membership inference to verify specific data participation, and statistical analysis to infer private attributes. While SSL encryption secures transmissions, it does not prevent adversaries from analyzing transmitted updates.
@@
To mitigate these risks, our approach integrates selective encryption, differential noise addition, and bitwise scrambling. Critical model parameters are encrypted, while unencrypted data is obfuscated with noise and scrambling to prevent reconstruction. Secure aggregation further ensures data protection throughout training, providing strong defenses against privacy attacks without incurring excessive computational overhead. This layered strategy effectively balances security and efficiency, making it well-suited for sensitive applications like healthcare.
IV Proposed Model
@@ IV-A Selective Encryption
IV-B Differential Noise Addition
@@ IV-C Bitwise Scrambling
IV-D Federated Learning Integration
@@
IV-E Security and Efficiency Trade-Off
@@
IV-F Privacy-Preserving Techniques
IV-F1 Homomorphic Encryption
@@
where Enc and Dec denote encryption and decryption, respectively, and Eval represents a computation on encrypted data. This property is critical for privacy-preserving applications, as it prevents adversaries from accessing plaintext data [16].
IV-F2 Differential Privacy
@@
where $\epsilon$ is the privacy budget, controlling the trade-off between privacy and utility [30].
@@
IV-G Process Flow of the Proposed Model
@@
IV-H Federated Learning System Process Flow
@@
IV-H1 Client-Side Processing
@@
IV-H2 Server-Side Processing
@@
IV-H3 Client Post-Processing
@@
IV-I Advantages of the Proposed Model
@@ IV-J1 Setup and Definitions
Parameters in $S$ are encrypted using a semantically secure HE scheme. Semantic security ensures no polynomial-time adversary can distinguish ciphertexts of different messages, implying no additional privacy cost in a DP sense (0-DP).[16]
@@
For each $i\in[N]\setminus S$, we add noise calibrated to $\epsilon_{i}$-DP. By the composition property of DP, releasing all these parameters together is $\left(\sum_{i\in[N]\setminus S}\epsilon_{i}\right)$-DP.[30]
@@
We define a scrambling function $\mathcal{T}_{k}:\mathcal{X}\to\mathcal{X}$, where $\mathcal{X}$ is the space of possible model outputs. $\mathcal{T}_{k}$ is deterministic, keyed by $k$, and independent of $D$ except through $M(D)$. This scrambling is a form of post-processing.[43]
IV-J2 Privacy Guarantee
Theorem 1.
@@ IV-J2 Privacy Guarantee
    Encrypts $W_{i}$ for $i\in S$ using a semantically secure HE scheme [16, 44].
@@
    Adds noise to each $W_{j}$ for $j\in[N]\setminus S$ to achieve $\epsilon_{j}$-DP. Releasing all noisy parameters together is $\left(\sum_{j\in[N]\setminus S}\epsilon_{j}\right)$-DP [42].
@@
    Applies a scrambling function $\mathcal{T}_{k}$ to the entire output, resulting in $\mathcal{M}^{\prime}(D)=\mathcal{T}_{k}(\mathcal{M}(D))$ [43].
@@
Proof.
Step 1 (DP on $[N]\setminus S$): For $[N]\setminus S$, each parameter $W_{j}+\text{noise}_{j}$ is $\epsilon_{j}$-DP. By composition, releasing all these parameters is $\sum_{j\in[N]\setminus S}\epsilon_{j}$-DP [42]. Formally, for any adjacent $D,D^{\prime}$ and event $O$:
@@
[42].
Step 2 (Selective Encryption on $S$): Parameters in $S$ are encrypted with a semantically secure HE scheme. Without the secret key, no information about the plaintext is revealed, contributing 0-DP cost [16, 44].
Step 3 (Scrambling as Post-Processing): The scrambling $\mathcal{T}_{k}$ is a deterministic, data-independent post-processing step. A fundamental property of DP is closure under post-processing: if $\mathcal{M}$ is $\epsilon$-DP, then $f(\mathcal{M})$ is also $\epsilon$-DP for any deterministic $f$ [43, 42]. Since $\mathcal{T}_{k}$ does not depend on $D$ beyond $\mathcal{M}(D)$ itself, applying $\mathcal{T}_{k}$ cannot increase or decrease the DP budget.
@@
Thus, $\mathcal{M}^{\prime}(D)$ maintains the same DP guarantee as $\mathcal{M}(D)$. ∎
V Evaluation
This section evaluates our federated learning setup using various security techniques, including homomorphic encryption, differential privacy, and our proposed FAS. We assess their impact on test accuracy, communication cost, computational overhead, and privacy preservation.
Each dataset was split into 10 equal partitions, distributed across machines running Flower clients. Each machine trained its local model over 10 rounds, with 20 epochs per round. A validation set (10% of training data) ensured consistency, and final accuracy was evaluated on a separate test set. All models were trained using CPU resources to simulate realistic constraints.
V-A Datasets Used in Experiments
We selected publicly available datasets covering diverse imaging contexts with high usability ratings. Our criteria included broad disease coverage, diverse imaging techniques, and sufficient training and validation data.
V-A1 CIFAR-10 Dataset
A benchmark for image classification, CIFAR-10 consists of 60,000 32x32 images across 10 classes. It includes 50,000 training and 10,000 test images, commonly used for model performance evaluation [46].
V-A2 Chest X-ray (COVID-19, Pneumonia) Dataset
This dataset contains 6,339 grayscale X-ray images across three classes: benign, COVID-19, and pneumonia (2,313 images per class). These images aid in lung condition assessment [7].